Show that Cp=Cv+R
1
Expert's answer
2020-09-15T13:53:43-0400

Consider one mole of ideal gas enclosed in a cylinder fitted with Mass less piston

According to first law of thermodynamics

dU=dq + dW

At constant volume ( isochoric)

dv=0

dU=dq

dq is also equal to nCvdT

[dU=nCvdT]

At constant pressure

dH=dU + PdV ,. dq=dU+PdV

dH=dq

[dH=nCpdT]

dH=dU+PdV

For an ideal gas equation

PdV=nRdT

dH=dU+nRdT

nCpdT=nCvdT+nRdT

Cp=Cv+R